# Prime and Reach: Synthesising Body Motion for Gaze-Primed Object Reach

### v1

@@ B.4 Train-Test Splits
Appendix C Ablation of Architecture & Loss
As in the main paper, our ablations are evaluated on HD-EPIC and MoGaze.
C.1 Transformer Encoder v/s Decoder
@@
We compare performance of transformer encoder v/s decoder based diffusion model for the task of P&R motion generation in Tab. S2. While the decoder architecture injects condition $\mathbf{\hat{z}_{t}}$ by cross-attention with each decoder layer, the encoder provides the condition as an additional token at the input of the first encoder layer. The decoder architecture performs significantly better than the encoder architecture, making it a superior choice for the task.
@@
Encoder 51.59 82.91 22.82 0.29 37.95 56.62 27.90 0.34
Decoder 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
C.2 Training Loss
We ablate the impact of $\mathcal{L}_{joint}$ in Tab. S3. We find adding the $\mathcal{L}_{joint}$ helps improve P&R generation for both HD-EPIC and MoGaze.
Table S3: Loss Ablation. The improvement by adding the joint loss is evident across both datasets and 7 out of 8 metrics.
@@ C.2 Training Loss
$\mathcal{L}$ 57.70 86.70 16.67 0.31 40.10 86.38 21.21 0.34
$\mathcal{L}+\mathcal{L}_{joint}$ 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
C.3 Incorporating Goal Condition
We pretrain our P&R diffusion model for motion generation only conditioned on text using the Nymeria dataset. For fine-tuning on P&R sequences, we add our initial state and goal pose/target location condition $\mathbf{p}$ to the text condition $\mathbf{z_{t}}$ to get $\mathbf{\hat{z}}_{t}$. We ablate another alternative of incorporating $\mathbf{p}$ to $\mathbf{z_{t}}$ using cross-attention as shown in
@@ C.3 Incorporating Goal Condition
where $CA$ is a 1-layer cross-attention. $\mathbf{z_{t}}$ is linearly projected to get the query and $\mathbf{p}$ is projected to key and value. We use a residual network to make the most of our pretraining. We show the results in Tab. S4. We find that incorporating the condition through addition performs better, across all metrics.
@@
CACA 56.41 85.67 17.33 0.27 39.43 72.62 26.04 0.35
Addition 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
C.4 Ablation of Total Diffusion Steps
For a single motion generation, the diffusion model starts from noise at $t=T$ and iteratively denoises it over diffusion steps $t=\{T,T-1,\cdots,0\}$, finally producing the clean motion at $t=0$. We ablate the choice of $T$ in Tab. S5, which controls the total number of steps needed to generate a sequence of motion. We find that $T=50$ gives a consistently good performance across all metrics with an improvement of $1.64\%$ in prime success. Importantly the method is generally robust to the number of steps.
@@ C.4 Ablation of Total Diffusion Steps
10 57.42 90.26 16.00 0.26 41.22 89.88 26.64 0.34
50 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
100 56.18 85.07 17.38 0.27 38.02 91.29 22.25 0.33
500 55.54 85.48 18.13 0.29 36.76 91.96 23.88 0.34
1000 54.68 85.22 17.67 0.29 37.65 90.25 26.04 0.34
Figure S3: Varying time window $\sigma$ and angular error threshold $\theta$ for Prime Success calculation on HD-EPIC and MoGaze.
Appendix D Analysing hyper-parameters of Prime Success metric
We conduct an in-depth analysis to better understand the impact of hyperparameters (the time window $\sigma$ and threshold for angular error $\theta$) used in calculating the newly introduced Prime Success metric. We compare our P&R predictions with the actual gaze while varying the hyperparameters of the metric. Note that as the thresholds are changed, the motion can be considered a success or a failure. Recall that our results are reported for $\theta=16\deg$ and $\sigma=0.2$ sec.
To evaluate the impact of these hyperparameters, we vary $\theta$ on the x-axis (between 0 and 90 $\deg$), then plot distinct curves for discrete time windows: 0, 0.2, 0.4, 0.8, and 1.0 seconds. Figure S3 shows that a very tight time window is too restrictive for MoGaze. As expected, a high $\theta$ threshold is too permissive and cannot be used to compare different methods.
Appendix E Pretraining Results
Here we provide the results of our text-conditioned motion generation pre-training on Nymeria.
As in previous works [tevet2023mdm], we train motion-text embedding models [Guo2022t2m], with two encoders: one for motion and one for text, using a contrastive loss. We use paired text-motion sequences from the Nymeria train set. We follow the architecture for our encoders from  [Guo2022t2m].
Following [guo2020action2motion, tevet2023mdm], we use the following metrics for evaluation -
@@ Appendix E Pretraining Results
E.1 Evaluating Embedding and Encoders
We first evaluate the quality of our trained encoders. We use the embedding from [tevet2023mdm] as our baseline, which is trained on HumaML3D. We compare the encoders on the matched text and real motion pairs of the Nymeria test set in Tab. S6.
Table S6: Evaluating encoders.
Feature
Extractor
R Precision
(Top- 3)
$\uparrow$
FID $\downarrow$
Multi-modal
Distance
$\downarrow$
Diversity $\uparrow$
[…]
Different from HumanML3D, Nymeria text and motion relate to more fine-grained everyday activities e.g., ‘In the hallway, C slightly turns her body to the left as she looks at her peer while holding the right side door of the laundry closet using her right hand. Then C slightly turns her body to the right as she looks at the laundry closet and takes her right hand off from the right side door of the laundry closet’. As a result, HumanML3D-trained text embeddings do not work well on the Nymeria test set. The high R-Precision and low multi-modal distance show that our trained embedding on Nymeria better matches text-motion pairs while maintaining high diversity. This shows that our trained embedding is substantially better and we thus use it for evaluating the impact of our pretraining.
E.2 Pretraining Results
We provide the results of our pretrained model in Tab. S7. We find that the motion generated by our Nymeria-pretrained model aligns better with the fine-grained texts of Nymeria. This is verified by the $34.0\%$ and 2.68-2.68 improvements in R-Precision and Multi-modal distance respectively. The diversity of our generated motions is +2.49+2.49 higher than that of motions generated by the HumanML3D pre-trained model. We showcase some of the qualitative results of the pre-trained model in Fig. S4.

### v2

@@ 0.B.5 Train-Test Splits
Appendix 0.C Per-Dataset Models
In the main paper, we train a single model on the combined training sequences of all datasets, and present our main results in Tab. 2. For completion, and to further validate the effectiveness of our proposed method, we evaluate the performance of the models trained and tested independently (i.e. separately) on each of the five datasets - training one model per dataset.
As shown in Tab.˜S3, the performance trends align closely with the results reported in the main manuscript. Even when trained on domain-specific data, the naive static and text-conditioned baselines perform poorly across all metrics. This confirms that the difficulty of the P&R task, specifically the need for spatial guidance and anticipatory motion, cannot be overcome by restricting training to a single dataset if the model architecture lacks sufficient conditioning cues. For conditioning with goal pose, our P&R model performs the highest prime success in all datasets and reach success in 4 out of 5 datasets. In particular, P&R outperforms the strongest baselines by $15.1\%$ and $19.9\%$ in prime success and reach success on the MoGaze dataset, respectively. A similar trend can be seen in the results with object location conditioning, where the proposed method achieves the highest prime success in most datasets while keeping reach success nearly perfect. Furthermore, our model consistently yields the lowest location error compared to all baselines. Taken together, these results demonstrate that the proposed method is robust across diverse domains and independent of the size of the training data.
Table S3: Per-Dataset Training - Comparison of motion generation baselines. Here, we train an independent model per dataset (HD-EPIC, MoGaze, HOT3D, ADT, and GIMO), and report results for each. The baselines are grouped by the type of conditioning used for generation. $\dagger$ denotes the zero-shot inference. For MDM, we evaluate two pre-trained models: (1) trained on HumanML3D [24], and (2) trained on Nymeria.[57], denoted as $\ddagger$ and , respectively. Entries without a marker correspond to models fine-tuned on our per-dataset P&R sequences.
HD-EPIC MoGaze
Condition Method
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
No condition Static 0 23.16 0.70 50.91 0.45 0 2.56 1.06 75.99 0.62
Text MDM $\dagger\ddagger$ 9.53 13.94 1.14 81.75 0.96 0.16 2.85 2.20 2.03 94.11 1.45 0.39
MDM $\dagger^{*}$ 9.52 13.47 0.85 62.13 0.59 0.06 3.11 1.85 1.19 79.76 0.73 0.06
MDM 14.30 19.85 0.82 52.26 0.51 0.06 6.81 4.50 1.19 82.62 0.74 0.04
+ Initial State & Goal Pose GMD $\dagger$[39] 33.27 30.77 0.26 2.00 0.32 0.10 8.11 10.86 0.35 2.23 0.54 0.11
GMD[39] 44.40 56.74 0.21 2.00 0.30 0.06 25.38 34.66 0.28 1.00 0.53 0.21
DNO $\dagger$[38] 43.42 87.44 0.07 2.60 0.30 0.04 17.39 69.71 0.09 6.81 0.64 0.07
DNO [38] 48.99 89.16 0.09 0.49 0.25 0.06 31.15 78.68 0.12 0.48 0.55 0.18
DartControl $\dagger$ [101] 30.06 81.23 0.14 2.20 0.43 0.15 38.26 71.16 0.11 0.14 0.52 0.35
DartControl [101] 34.65 88.46 0.11 1.78 0.37 0.09 34.78 65.94 0.12 0.14 0.58 0.51
P&R 53.45 95.60 0.08 0.49 0.20 0.05 53.33 98.55 0.08 0.00 0.49 0.17
+ Initial State & Object Loc. WANDR $\dagger$[15] 33.28 80.92 0.47 40.35 0.42 0.11 31.74 96.81 0.62 60.14 0.61 0.25
WANDR [15] 33.54 74.54 0.54 47.91 0.49 0.16 49.42 98.70 0.69 68.99 0.66 0.24
DNO $\dagger$[38] 37.34 100.00 0.47 37.87 0.44 0.05 27.24 100.00 0.64 59.56 0.87 0.09
DNO [38] 46.10 100.00 0.42 33.33 0.29 0.05 32.40 100.00 0.52 45.43 0.71 0.16
DartControl $\dagger$ [101] 28.82 89.20 0.47 40.82 0.52 0.13 42.90 100.00 0.54 50.29 0.72 0.40
DartControl [101] 30.68 89.29 0.44 35.56 0.46 0.08 42.03 100.00 0.56 52.17 0.75 0.54
P&R 53.70 100.00 0.37 24.36 0.25 0.04 61.59 100.00 0.51 42.44 0.69 0.17
HOT3D ADT
Condition Method
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
No condition Static 0 26.43 0.35 22.01 0.32 0 9.90 1.86 69.27 1.03
Text MDM $\dagger\ddagger$ 27.28 6.25 1.11 75.85 0.89 0.31 9.38 5.21 2.54 97.40 1.76 0.35
MDM $\dagger^{*}$ 8.65 4.30 0.51 36.20 0.44 0.00 3.65 6.25 1.98 83.33 1.16 0.06
MDM 27.86 34.65 0.45 18.45 0.37 0.00 11.76 17.64 1.95 89.41 1.20 0.18
+ Initial State & Goal Pose GMD $\dagger$[39] 31.85 20.16 0.38 25.00 0.37 0.02 16.47 12.94 0.35 10.58 0.48 0.21
GMD [39] 46.84 73.39 0.21 6.87 0.34 0.03 22.35 25.88 0.32 9.41 0.43 0.22
DNO $\dagger$[38] 45.83 90.18 0.05 9.67 0.23 0.02 25.88 52.94 0.04 5.88 0.56 0.08
DNO [38] 47.10 92.11 0.03 0.72 0.18 0.02 28.23 82.35 0.05 1.17 0.43 0.16
DartControl $\dagger$ [101] 40.51 89.77 0.17 2.15 0.24 0.01 14.12 67.06 0.22 8.24 0.62 0.24
DartControl [101] 46.84 89.91 0.17 2.15 0.27 0.00 23.53 56.47 0.33 10.59 0.68 0.19
P&R 53.22 93.81 0.06 0.00 0.11 0.00 37.64 70.58 0.08 1.17 0.33 0.11
+ Initial State & Object Loc. WANDR $\dagger$[15] 34.32 91.66 0.33 17.77 0.27 0.05 7.65 82.94 0.66 61.76 0.60 0.22
WANDR [15] 41.45 91.66 0.33 17.50 0.24 0.04 38.24 80.00 0.66 58.24 0.57 0.21
DNO $\dagger$ [38] 46.77 100.00 0.41 40.86 0.36 0.04 23.53 100.00 0.64 61.18 0.70 0.10
DNO [38] 58.06 100.00 0.35 20.16 0.30 0.01 38.82 100.00 0.60 57.79 0.59 0.10
DartControl $\dagger$ [101] 46.43 97.58 0.38 19.78 0.34 0.01 25.88 97.65 0.69 64.71 0.86 0.29
DartControl [101] 45.09 96.64 0.33 9.15 0.30 0.00 31.76 96.47 0.69 72.94 0.75 0.21
P&R 65.45 100.00 0.25 4.16 0.16 0.01 43.53 100.00 0.55 52.54 0.48 0.14
GIMO
Condition Method
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
No condition Static 0 0 3.41 100.0 1.86
Text MDM $\dagger\ddagger$ 0 0 3.69 100 2.17 0.45
MDM $\dagger^{*}$ 0 0 3.34 100 1.82 0.08
MDM 0 0 2.74 100 1.56 0.12
+ Initial State & Goal Pose GMD $\dagger$[39] 13.63 4.76 0.44 11.90 0.62 0.43
GMD [39] 22.72 9.09 0.34 9.09 0.58 0.19
DNO $\dagger$ [38] 18.18 27.27 0.15 9.09 0.80 0.18
DNO [38] 31.82 36.36 0.16 4.54 0.62 0.18
DartControl $\dagger$ [101] 9.52 14.29 0.33 9.52 1.37 0.07
DartControl [101] 38.10 9.52 0.35 9.52 0.85 0.04
P&R 45.45 59.09 0.13 4.54 0.55 0.18
+ Initial State & Object Loc. WANDR $\dagger$[15] 14.29 61.90 0.51 57.14 0.79 0.44
WANDR [15] 14.29 66.67 0.51 47.62 0.57 0.43
DNO $\dagger$ [38] 27.27 100.00 0.76 63.63 0.98 0.11
DNO [38] 36.36 100.00 0.45 50.00 0.70 0.12
DartControl $\dagger$ [101] 4.76 76.19 0.49 61.90 1.51 0.10
DartControl [101] 57.14 66.67 0.44 47.62 0.87 0.05
P&R 50.00 100.00 0.38 36.36 0.63 0.18
Appendix 0.D Ablation of Architecture & Loss
We present additional ablations of our architecture and losses. As in the main paper, we train a single model on all datasets, and report results on two datasets: HD-EPIC and MoGaze.
0.D.1 Transformer Encoder v/s Decoder
@@
We compare performance of transformer encoder v/s decoder based diffusion model for the task of P&R motion generation in Tab.˜S4. While the decoder architecture injects condition $\mathbf{\hat{z}_{t}}$ by cross-attention with each decoder layer, the encoder provides the condition as an additional token at the input of the first encoder layer. The decoder architecture performs significantly better than the encoder architecture, making it a superior choice for the task.
@@
Encoder 35.18 94.05 56.19 0.54 56.80 90.89 59.29 0.85
Decoder 51.00 100.00 25.26 0.27 62.37 100.00 36.43 0.56
0.D.2 Training Loss
We ablate the impact of $\mathcal{L}_{\text{joint}}$ in Tab.˜S5. We find adding the $\mathcal{L}_{\text{joint}}$ helps improve P&R generation for both HD-EPIC and MoGaze.
Table S5: Loss Ablation.
@@ 0.D.2 Training Loss
$\mathcal{L}$ 51.65 100.00 33.22 0.34 61.10 100.00 57.24 0.70
$\mathcal{L}+\mathcal{L}_{\text{joint}}$ 51.00 100.00 25.26 0.27 62.37 100.00 36.43 0.56
0.D.3 Incorporating Goal Condition
[…]
@@ 0.D.3 Incorporating Goal Condition
where $CA$ is a 1-layer cross-attention. $\mathbf{z_{t}}$ is linearly projected to get the query and $\mathbf{p}$ is projected to key and value. We use a residual network to make the most of our pretraining. We show the results in Tab.˜S6. We find that incorporating the condition through addition performs better, across 7 out of the 8 metrics.
@@
CACA 50.20 100.00 33.45 0.36 65.30 100.00 49.34 0.68
Addition 51.00 100.00 25.26 0.27 62.37 100.00 36.43 0.56
0.D.4 Ablation of Number of Diffusion Steps $T$
For a single motion generation, the diffusion model starts from noise at $t=T$ and iteratively denoises it over diffusion steps $t=\{T,T-1,\cdots,0\}$, finally producing the clean motion at $t=0$. We ablate the choice of $T$ in Tab.˜S7, which controls the total number of steps needed to generate a sequence of motion. We find that $T=50$ gives consistently good performance across all metrics, with at least a $1.30\%$ improvement in prime success on HD-EPIC. Importantly, the method is generally robust to the number of steps.
@@ 0.D.4 Ablation of Number of Diffusion Steps $T$
10 49.45 100.00 28.46 0.30 63.50 100.00 54.02 0.64
50 51.00 100.00 25.26 0.27 62.37 100.00 36.43 0.56
100 50.30 100.00 25.76 0.28 62.00 100.00 48.86 0.61
500 49.80 100.00 27.46 0.31 61.00 100.00 49.95 0.65
1000 49.65 100.00 27.01 0.33 61.45 100.00 51.10 0.68
Figure S3: Varying time window $\sigma$ and proximity threshold $\delta$ for Prime Success calculation on HD-EPIC and MoGaze.
Appendix 0.E Prime Success Metric: Analysis
We conduct an in-depth analysis to better understand the impact of hyperparameters (the time window $\sigma$ and proximity threshold $\delta$ for determining gaze-object intersection) used in calculating the newly introduced Prime Success metric. We compare our P&R predictions while varying the hyperparameters of the metric. Note that as the thresholds are changed, the motion can be considered a success or a failure. Recall that our results are reported for $\delta=0.25m$ and $\sigma=1.0$ sec.
To evaluate the impact of these hyperparameters, we vary $\delta$ on the x-axis (between 0 and 1 m), then plot distinct curves for discrete time windows: 0, 0.25, 0.5, 0.75, 1.0, 1.25 seconds. Fig.˜S3 shows that a very tight time window is too restrictive for MoGaze. As expected, a high $\delta$ threshold is too permissive and cannot be used to compare different methods.
Appendix 0.F Evaluating Pre-trained Models
In Tab. 4 in the main manuscript, we presented results showcasing the impact of the pre-training dataset on our model’s performance. For completion, we here also present results on evaluating the pre-trained models themselves, before we do any fine-tuning. We evaluate the models trained on HumanML3D and Nymeria for the task of text-conditioned motion generation.
As in previous works [84], we train motion-text embedding models [23], with two encoders: one for motion and one for text, using a contrastive loss. We use paired text-motion sequences from the Nymeria train set. We follow the architecture for our encoders from  [23].
Following [26, 84], we use the following metrics for evaluation -
@@ Appendix 0.F Evaluating Pre-trained Models
We provide the results of our pretrained model in Tab.˜S8. We find that the motion generated by our Nymeria-pretrained model aligns better with the fine-grained texts of Nymeria. This is verified by the $34.0\%$ and 2.68-2.68 improvements in R-Precision and Multi-modal distance respectively. The diversity of our generated motions is +2.49+2.49 higher than that of motions generated by the HumanML3D pre-trained model. We showcase some of the qualitative results of the pre-trained model in Fig.˜S4.